Find the probability of randomly selecting 2 mathematics books and 3 physics books from a box containing 4 mathematics books and 8 physics books.
Answer by stanbon(57984)

Find the probability of randomly selecting 2 mathematics books and 3 physics books from a box containing 4 mathematics books and 8 physics books.
----
ways to succeed: [4C2*8C3] = 336
possible outcomes: 12C5 = 792
---
Prob = 336/792 = 0.4242..
Cheers,
Stan H.